OpenStudy (anonymous):

PLEASE HELP ME! The value of 3^1003 +3^1002 divided by 3^1001 - 3^1000 is: ?

OpenStudy (anonymous):

Explain your answer please!

OpenStudy (anonymous):

When you multiply two numbers with powers the powers are additive provide both numbers have the same base. As an example (a^4)(a^6)=a^10............a is the base and 4+6 which are the powers add to 10. If you divide two numbers with powers, then we subtract the power of the numerator from the denominator provide both numbers have the same base. ie a^6/a^4=a^2. As a hint to your problem pull out the highest power from 1003^3 + 1002^3 and then pull out the highest number from 1001^3 -1000^3. You want to write in the following form 1000^?(??^3 + ???^3) so ?, ??. and ??? represent three different numbers. I hope this helps.

OpenStudy (anonymous):

um i don't really understand what you are saying.... this is all confusing me. would u help me solve this specific problem?

OpenStudy (anonymous):

\[\frac{3{}^{\wedge}1003+3{}^{\wedge}1002 }{3{}^{\wedge}1001-3{}^{\wedge}1000} \]\[\frac{3{}^{\wedge}3 *3{}^{\wedge}1000+3{}^{\wedge}2*3{}^{\wedge}1000}{3{}^{\wedge}1*3{}^{\wedge}1000-3{}^{\wedge}0*3{}^{\wedge}1000} \]\[\frac{3{}^{\wedge}3 *a+3{}^{\wedge}2*a}{3{}^{\wedge}1*a-3{}^{\wedge}0*a}\text{=}\frac{(36 a)}{(2 a)}\text{=}18\]

OpenStudy (anonymous):

thank you so much! that helped a lot!

OpenStudy (anonymous):

"a" is an arbitrary symbol replacing the expression 3^1000. It eventually divides out. Thank you for the medal.